OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (-1, -3) and (2, 1)

OpenStudy (***[isuru]***):

first u can write the standard for for an line.. y = mx + b where m = slope b = intercept to find slope... if u have 2 sets of coordinates.. (x1 , y) and (x2 , y2) \[slope = \frac{ y_{2} - y_{1} }{ x_{2} - x_{2} }\] since u got a pair of coordinates what will be ur slope ?

OpenStudy (afifarahman):

Since two points are given, so two point formula will be used and equation will come out to be \[4x-3y-5=0\]

OpenStudy (anonymous):

i only have 7 spaces to put the standard form in for example. (x3+x5=8)

OpenStudy (***[isuru]***):

if u simplify it ... u will gain a slope of ... 4/3 which means u will have an equation of y= (4/3)x + b now to find b lets use (2 , 1) AND plug the in to our equation.. 1 = (4/3)2 + b b= -5/3.. so... ur equation will be y= (4/3)x - 5/3 which is 3y = 4x -5 got it :) ?

OpenStudy (anonymous):

thanks

OpenStudy (***[isuru]***):

u r welcome!